# N. Deng & Y. Wang & W. Cao

### v1

@@ Abstract
This paper investigates the stochastic Cahn–Hilliard equation (SCHE) driven by additive space–time white noise. We first refine the analytical ergodic theory by proving that the continuum equation admits a unique invariant measure in the more regular state space $H_{\alpha}$, extending the classical result of Da Prato & Debussche (1996) on the negative Sobolev space $\dot{H}^{-1}_{\alpha}$. To approximate long-time behaviour, we introduce an explicit fully discrete scheme that combines a finite-difference spatial discretization with a strongly tamed exponential Euler method in time. Uniform-in-time moment bounds in the $L^{\infty}$-norm are established for the numerical solution, and a uniform strong convergence estimate with an explicit rate is derived for the fully discrete approximation. Exploiting a mass-preserving minorization tailored to Neumann boundary conditions, we further show that the numerical scheme is geometrically ergodic and possesses a unique invariant measure, together with polynomial-order error bounds for approximating the exact invariant measure. Strong laws of large numbers are proved for both the continuous and discrete systems, ensuring almost-sure convergence of temporal averages to the corresponding ergodic limits. Numerical experiments corroborate the theoretical findings, including the long-time strong convergence and the accuracy of invariant measure approximation. Overall, the results provide a complete analytical and numerical framework for investigating the long-time statistical behaviour of the SCHE. Stochastic Cahn–Hilliard equation; ergodicity; fully discrete schemes; approximation of invariant measures; law of large numbers
@@ 1 Introduction
$\frac{\partial u}{\partial t}+\Delta^{2}u=\Delta f(u)+\sigma\frac{\partial W}{\partial t},\quad\text{in }[0,\infty)\times\mathcal{O},$ (1a)
@@
$\frac{\partial u}{\partial n}=\frac{\partial\Delta u}{\partial n}=0,\qquad\qquad\qquad\quad\text{on }[0,\infty)\times\partial\mathcal{O},$ (1b)
$u(0,\cdot)=u_{0},$ (1c)
@@
with a cubic polynomial $f$ and additive space–time white noise is a prototypical phase-field model for binary alloys subject to thermal fluctuations (Cahn, 1961; Cook, 1970; Langer, 1971). Its fourth-order dissipative structure, non-globally Lipschitz drift, and mass conservation under Neumann boundary conditions jointly create substantial analytical and numerical challenges, particularly in the study of long-time behaviour and invariant measures.
@@
The ergodic properties of the SCHE have been established only in low-regularity spaces. The classical work of Da Prato & Debussche (1996) proved existence and uniqueness of an invariant measure on the negative Sobolev space $\dot{H}^{-1}_{\alpha}$, and subsequent refinements for singular nonlinearities or degenerate noise (Debussche & Zambotti, 2007; Debussche & Goudenège, 2011; Goudenège & Manca, 2015) remain confined to this setting. The only existence result in a more regular space—due to Jiang et al. (2014)—does not yield uniqueness. Consequently, the ergodic theory of the SCHE in physically relevant Hilbert spaces remains incomplete.
@@
On the numerical side, substantial progress has been made on finite-time discretizations of the SCHE. Strong convergence rates have been obtained for additive noise using Galerkin finite element or spectral methods combined with backward Euler, accelerated implicit Euler, or tamed exponential Euler schemes (Qi & Wang, 2020; Cui et al., 2021; Cai et al., 2023a), and for fractional Brownian forcing via finite difference and tamed exponential integrators (Deng& Cao, 2025). Related advances for multiplicative noise appear in (Cui & Hong, 2020, 2023; Hong et al., 2024). However, all these works are restricted to finite-time analysis, and none address uniform moment bounds, ergodicity, or invariant measure approximation.
@@
Invariant measure approximation for SPDEs is well developed for linear problems and semilinear equations with globally Lipschitz or monotone drifts (Bréhier & Vilmart, 2016; Chen et al., 2017, 2020; Hong & Wang, 2019; Cui et al., 2021). These works typically rely on implicit time discretizations, whose computational cost becomes prohibitive for long-time simulations. Motivated by this, recent studies have explored explicit exponential or tamed exponential schemes for nonlinear SPDEs, yielding polynomial-in-time or uniform moment bounds for the stochastic Allen–Cahn and Burgers–Huxley equations (Bréhier, 2022; Wang et al., 2025; Wang & Cao, 2024). The unified explicit framework of Jiang & Wang (2025) further demonstrates that geometric ergodicity can be achieved for a broad class of semilinear SPDEs. Nevertheless, none of these approaches apply to the SCHE: the cubic nonlinearity appears under an additional Laplacian, which leads to an unbounded operator applied to a non-globally Lipschitz drift; combined with the fourth-order structure and the mass constraint imposed by the Neumann boundary condition, this destroys the dissipativity and minorization arguments used in existing long-time analyses for second-order equations such as the stochastic Allen–Cahn equation. Consequently, long-time accuracy, ergodicity, and invariant measure approximation for the SCHE remain completely open.
@@
Second, we construct an explicit fully discrete scheme by combining a finite difference spatial discretization with a strongly tamed exponential Euler method in time. The taming strategy, originally introduced in Wang & Cao (2024), is incorporated into our fourth-order setting and allows us to establish uniform-in-time moment bounds for the fully discrete solution—a property that has not previously been proved for numerical approximations of the SCHE. These bounds form the key ingredient for our long-time analysis. On this basis, we prove a uniform strong convergence result
@@
Third, we develop a new ergodicity framework suitable for Neumann boundary conditions. Since the conserved mass invalidates the classical minorization condition of Mattingly, Stuart & Higham (2002), we introduce a mass-preserving minorization posed on affine hyperplanes and combine it with a coupling argument to prove existence, uniqueness, and geometric ergodicity of the numerical invariant measure. Furthermore, we obtain a quantitative approximation estimate
@@ 2 Preliminaries
Considering the Neumann boundary condition, we define the subspace $\dot{H}:=\left\{v\in H:\left.\frac{\partial v}{\partial n}\right|_{x\in\partial O}=0\right\}$. The operator $A:=\Delta$ denotes the Laplacian with Neumann boundary conditions, whose domain is $\mathcal{D}(A):=H^{2}\cap\dot{H}$. There exists an orthonormal basis $\left\{\phi_{j}\right\}_{j=0}^{\infty}$ of $\dot{H}$ and an increasing sequence of eigenvalues $\left\{\lambda_{j}\right\}_{j=0}^{\infty}$ satisfying $A\phi_{j}=\lambda_{j}\phi_{j}$. Specifically, $\phi_{0}(x)=\sqrt{1/\pi}$ and $\phi_{j}(x)=\sqrt{2/\pi}\cos(jx)$ for $x\in\mathcal{O}$ and $j\geq 1$. The corresponding eigenvalues are $\lambda_{j}=j^{2}$ for $j\geq 0$. We define the fractional powers of A-A on $\dot{H}$ by
@@
To study the invariant measure, we follow Da Prato & Debussche (1996) and introduce the spaces corresponding to $\alpha\in\mathbb{R}$:
$\dot{H}^{-1}_{\alpha}:=\left\{v\in\dot{H}^{-1}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}\ \text{and}\ H_{\alpha}:=\left\{v\in\dot{H}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}.$ (2)
@@
$e^{-x}\leq Cx^{-\alpha}\quad\text{and}\quad 1-e^{-x}\leq Cx^{\hat{\alpha}},\quad x>0.$ (3)
@@
$\left\|(-A)^{\gamma}e^{-A^{2}t}\hat{u}\right\|_{L^{2}}\leq C\left(t^{-\frac{\gamma}{2}}e^{-\frac{\lambda_{1}^{2}}{2}t}\left\|\hat{u}\right\|_{L^{2}}+g(\gamma)\left|\left\langle\hat{u},\phi_{0}\right\rangle_{L^{2}}\right|\right),\quad t>0,\$(4)
$\left\|(-A)^{\gamma_{1}}e^{-A^{2}t}\left(I-e^{-A^{2}s}\right)\hat{u}\right\|_{L^{2}}\leq C\left(t^{-\frac{\gamma_{1}-\gamma_{2}+2\gamma_{3}}{2}}s^{\gamma_{3}}e^{-\frac{\lambda_{1}^{2}}{2}t}\left\|(-A)^{\gamma_{2}}\hat{u}\right\|_{L^{2}}+g(\gamma_{1})\left|\left\langle\hat{u},\phi_{0}\right\rangle_{L^{2}}\right|\right),\quad s,t>0,$ (5)
@@
We next state the assumptions on the nonlinear term $f$, the initial value $u_{0}$, and the noise $W(t,x)$. {assumption} Suppose that $f$ satisfies the following two conditions:
@@
(ii) $L_{f}<\lambda_{1}$, where $L_{f}=-\sup_{x\in\mathcal{O}}f^{\prime}(x)$. The above assumption ensures that $f$ satisfies
$(y-x)\bigl(f(x)-f(y)\bigr)$ $\leq L_{f}\,|x-y|^{2},\qquad\ \qquad\quad\ x,y\in\mathbb{R},$ (6)
$|f(x)-f(y)|$ $\leq C_{f}\bigl(1+x^{2}+y^{2}\bigr)|x-y|,\quad x,y\in\mathbb{R},$ (7)
$(1-\epsilon_{0})\left\langle A\hat{u},\hat{u}\right\rangle_{L^{2}}+\left\langle F(\hat{u}-\hat{v})-F(\hat{v}),\hat{u}\right\rangle_{L^{2}}$ $\leq-\epsilon_{0}\|\hat{u}\|_{L^{2}},\qquad\qquad\ \qquad\hat{u},\hat{v}\in L^{6}(\mathcal{O}),$ (8)
@@
{assumption}
Suppose that the initial condition $u_{0}\in H^{1}.$ {assumption} Suppose that the noise $W(t,\cdot)$ is a cylindrical Wiener process on $\dot{H}$, which can be represented by the formal series
@@
The concepts of the invariant measure and the Feller property are introduced below, following Hong & Wang (2019) for details. For a given Banach space $\mathcal{X}$, we denote by $B_{b}(\mathcal{X})$ (resp. $C_{b}(\mathcal{X})$) the Banach space of all bounded, Borel-measurable functions (resp. bounded continuous functions) on $\mathcal{X}$. Furthermore, let $\mathcal{L}(\mathcal{X})$ denote the space of all bounded linear operators on $\mathcal{X}$. We define the transition semigroup $P_{t}:[0,\infty)\rightarrow\mathcal{L}\left(B_{b}(\mathcal{X})\right)$ by
@@
where $u(u_{0};t,\cdot)$ is the solution of the SCHE (1) with initial condition $u(0,\cdot)=u_{0}(\cdot)$. It can be verified that $\left\{P_{t}\right\}_{t\geq 0}$ is a Markov semigroup.
@@
In Definition 2.4, we define the invariant measure for the transition semigroup $P_{t}$, which is uniquely determined by the SCHE (1). Therefore, in the following, the invariant measure for the semigroup, the SCHE, or the solution will refer to the same object and be used interchangeably.
3 Invariant measure for the stochastic Cahn–Hilliard equation
This section is devoted to the investigation of the invariant measure for the SCHE (1). In Da Prato & Debussche (1996), the authors established the existence and uniqueness of the invariant measure for (1) on the space $\dot{H}^{-1}_{\alpha}$. In this work, we extend these results to the space $H_{\alpha}$. Employing similar arguments to those in Da Prato & Debussche (1996), we directly conclude that the SCHE (1) admits a unique mild solution $u$ given by
$u(t,x)=$ $\int_{\mathcal{O}}G_{t}(x,y)u_{0}\left(y\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta G_{t-s}(x,y)f\left(u\left(s,y\right)\right)\mathrm{d}y\mathrm{d}s+\sigma\sum_{j=1}^{\infty}\int_{0}^{t}\int_{\mathcal{O}}G_{t-s}(x,y)\phi_{j}(y)\mathrm{d}y\mathrm{d}\beta_{j}(s),$ (9)
@@ 3 Invariant measure for the stochastic Cahn–Hilliard equation
$G_{t}(x,y):=\sum_{j=0}^{\infty}e^{-\lambda_{j}^{2}t}\phi_{j}(x)\phi_{j}(y),\quad t\in[0,\infty),\ x,\ y\in\mathcal{O}.$ (10)
@@
$\int_{\mathcal{O}}|G_{t}(x,y)-G_{t}(z,y)|^{2}\mathrm{d}y\leq e^{-\lambda_{1}t}\sum_{j=0}^{\infty}\lambda_{j}e^{-\lambda_{j}^{2}t}\left|x-z\right|^{2},\quad t\geq 0,\ x,z\in\mathcal{O},$ (11)
$\int_{\mathcal{O}}|\Delta G_{t}(x,y)-\Delta G_{t}(z,y)|\mathrm{d}y\leq Ce^{-\frac{\lambda_{1}}{2}t}\left|x-z\right|^{\alpha}\sqrt{\sum_{j=0}^{\infty}\lambda_{j}^{2+\alpha}e^{-\lambda_{j}^{2}t}},\quad t\geq 0,\ x,z\in\mathcal{O}.$ (12)
For the subsequent analysis, we rewrite (1) as
$\left\{\begin{aligned} &\mathrm{d}U_{t}+A^{2}U_{t}\mathrm{d}t=AF\left(U_{t}\right)\mathrm{d}t+\sigma\mathrm{d}W_{t},\quad t\in(0,\infty),\\ &U_{0}=u_{0}\in H.\end{aligned}\right$. (13)
Eq. (13) possesses a unique mild solution given by
@@
3.1 Key estimates for two auxiliary equations
@@
$\left\{\begin{aligned} &\mathrm{d}V_{t}+A^{2}V_{t}\mathrm{d}t=AF\left(V_{t}+\mathscr{O}_{t}\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &V_{0}=u_{0}.\end{aligned}\right$. (14)
@@
$\left\|v\right\|_{L^{6}}\leq C\left(\left\|v\right\|_{L^{2}}^{2}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}^{2}+\left\|Av\right\|_{L^{2}}^{2}\right)^{\frac{1}{12}}\left\|v\right\|_{L^{2}}^{\frac{5}{6}}\leq C\left(\left\|Av\right\|_{L^{2}}^{\frac{1}{6}}\left\|v\right\|_{L^{2}}^{\frac{5}{6}}+\left\|v\right\|_{L^{2}}\right),\quad v\in H^{2},$ (15)
@@
$\left\|v\right\|_{L^{\infty}}\leq C\left(\left\|v\right\|_{L^{2}}^{2}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}^{2}\right)^{\frac{1}{4}}\left\|v\right\|_{L^{2}}^{\frac{1}{2}}\leq C\left(\left\|v\right\|_{L^{2}}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}\right),\quad v\in H^{1}.$ (16)
In the next proposition, we present a general result. In particular, when $Z_{t}=\mathscr{O}_{t}$ and $y_{0}=u_{0}$ (17), we obtain the estimate for (14).
@@
$\left\{\begin{aligned} &\mathrm{d}Y_{t}+A^{2}Y_{t}\mathrm{d}t=AF\left(Y_{t}+Z_{t}\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &Y_{0}=y_{0},\end{aligned}\right$. (17)
@@
$\|Y_{t}\|^{2}_{L^{2}}+\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\|AY_{s}\|^{2}_{L^{2}}\mathrm{d}s\leq C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{2}+K_{1}(t)\right).$ (19)
Taking the inner product of (17) with $Y_{t}$ yields
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|Y_{t}\right\|^{2}_{L^{2}}\right)=$ $-\left\|AY_{t}\right\|^{2}_{L^{2}}+\left\langle F(Y_{t}+Z_{t}),AY_{t}\right\rangle_{L^{2}}.$ (20)
@@
$\left\langle F(Y_{t}+Z_{t}),AY_{t}\right\rangle_{L^{2}}=$ $\int_{\mathcal{O}}\left(a_{0}\left(y(t,x)+z(t,x)\right)^{3}+a_{1}\left(y(t,x)+z(t,x)\right)^{2}+a_{2}\left(y(t,x)+z(t,x)\right)+a_{3}\right)\partial_{xx}y(t,x)\mathrm{d}x$

### v3

# Ergodicity and invariant measure approximation of the stochastic Cahn–Hilliard equation via an explicit fully discrete scheme

@@ Abstract
This paper investigates the stochastic Cahn–Hilliard equation (SCHE) driven by additive space–time white noise. We first refine the analytical ergodic theory by proving that the continuum equation admits a unique invariant measure in the more regular state space $H_{\alpha}$, extending the classical result of [15] on the negative Sobolev space $\dot{H}^{-1}_{\alpha}$. To approximate long-time behaviour, we introduce an explicit fully discrete scheme that combines a finite-difference spatial discretization with a strongly tamed exponential Euler method in time. Uniform-in-time moment bounds in the $L^{\infty}$-norm are established for the numerical solution, and a uniform strong convergence estimate with an explicit rate is derived for the fully discrete approximation. Exploiting a mass-preserving minorization tailored to Neumann boundary conditions, we further show that the numerical scheme is geometrically ergodic and possesses a unique invariant measure, together with polynomial-order error bounds for approximating the exact invariant measure. Strong laws of large numbers are proved for both the continuous and discrete systems, ensuring almost-sure convergence of temporal averages to the corresponding ergodic limits. Numerical experiments corroborate the theoretical findings, including the long-time strong convergence and the accuracy of invariant measure approximation. Overall, the results provide a complete analytical and numerical framework for investigating the long-time statistical behaviour of the SCHE.
Keywords: Stochastic Cahn–Hilliard equation; ergodicity; fully discrete schemes; approximation of invariant measures; law of large numbers
@@ 1.  Introduction
$\frac{\partial u}{\partial t}+\Delta^{2}u=\Delta f(u)+\sigma\frac{\partial W}{\partial t},\quad\text{in }[0,\infty)\times\mathcal{O},$ (1.1a)
@@
$\frac{\partial u}{\partial n}=\frac{\partial\Delta u}{\partial n}=0,\qquad\qquad\qquad\quad\text{on }[0,\infty)\times\partial\mathcal{O},$ (1.1b)
$u(0,\cdot)=u_{0},$ (1.1c)
@@
with a cubic polynomial $f$ and additive space–time white noise is a prototypical phase-field model for binary alloys subject to thermal fluctuations [4, 11, 26]. Its fourth-order dissipative structure, non-globally Lipschitz drift, and mass conservation under Neumann boundary conditions jointly create substantial analytical and numerical challenges, particularly in the study of long-time behaviour and invariant measures.
@@
The ergodic properties of the SCHE have been established only in low-regularity spaces. The classical work of [15] proved existence and uniqueness of an invariant measure on the negative Sobolev space $\dot{H}^{-1}_{\alpha}$, and subsequent refinements for singular nonlinearities or degenerate noise [17, 18, 22] remain confined to this setting. The only existence result in a more regular space—due to [29]—does not yield uniqueness. Consequently, the ergodic theory of the SCHE in physically relevant Hilbert spaces remains incomplete.
@@
On the numerical side, substantial progress has been made on finite-time discretizations of the SCHE. Strong convergence rates have been obtained for additive noise using Galerkin finite element or spectral methods combined with backward Euler, accelerated implicit Euler, or tamed exponential Euler schemes [33, 13, 5], and for fractional Brownian forcing via finite difference and tamed exponential integrators [19]. Related advances for multiplicative noise appear in [12, 14, 23]. However, all these works are restricted to finite-time analysis, and none address uniform moment bounds, ergodicity, or invariant measure approximation.
@@
Invariant measure approximation for SPDEs is well developed for linear problems and semilinear equations with globally Lipschitz or monotone drifts [2, 9, 10, 24, 13]. These works typically rely on implicit time discretizations, whose computational cost becomes prohibitive for long-time simulations. Motivated by this, recent studies have explored explicit exponential or tamed exponential schemes for nonlinear SPDEs, yielding polynomial-in-time or uniform moment bounds for the stochastic Allen–Cahn and Burgers–Huxley equations [3, 36, 35]. The unified explicit framework of [25] further demonstrates that geometric ergodicity can be achieved for a broad class of semilinear SPDEs. Nevertheless, none of these approaches apply to the SCHE: the cubic nonlinearity appears under an additional Laplacian, which leads to an unbounded operator applied to a non-globally Lipschitz drift; combined with the fourth-order structure and the mass constraint imposed by the Neumann boundary condition, this destroys the dissipativity and minorization arguments used in existing long-time analyses for second-order equations such as the stochastic Allen–Cahn equation. Consequently, long-time accuracy, ergodicity, and invariant measure approximation for the SCHE remain completely open.
@@
Second, we construct an explicit fully discrete scheme by combining a finite difference spatial discretization with a strongly tamed exponential Euler method in time. The taming strategy, originally introduced in [35], is incorporated into our fourth-order setting and allows us to establish uniform-in-time moment bounds for the fully discrete solution—a property that has not previously been proved for numerical approximations of the SCHE. These bounds form the key ingredient for our long-time analysis. On this basis, we prove a uniform strong convergence result
@@
Third, we develop a new ergodicity framework suitable for Neumann boundary conditions. Since the conserved mass invalidates the classical minorization condition of [31], we introduce a mass-preserving minorization posed on affine hyperplanes and combine it with a coupling argument to prove existence, uniqueness, and geometric ergodicity of the numerical invariant measure. Furthermore, we obtain a quantitative approximation estimate
@@ 2.  Preliminaries
Considering the Neumann boundary condition, we define $\dot{H}:=\left\{v\in H:\left.\frac{\partial v}{\partial n}\right|_{x\in\partial O}=0\right\}$. The operator $A:=\Delta$ denotes the Laplacian with Neumann boundary conditions, whose domain is $\mathcal{D}(A):=H^{2}\cap\dot{H}$. There exists an orthonormal basis $\left\{\phi_{j}\right\}_{j=0}^{\infty}$ of $\dot{H}$ and an increasing sequence of eigenvalues $\left\{\lambda_{j}\right\}_{j=0}^{\infty}$ satisfying $A\phi_{j}=\lambda_{j}\phi_{j}$. Specifically, $\phi_{0}(x)=\sqrt{1/\pi}$ and $\phi_{j}(x)=\sqrt{2/\pi}\cos(jx)$ for $x\in\mathcal{O}$ and $j\geq 1$. The corresponding eigenvalues are $\lambda_{j}=j^{2}$ for $j\geq 0$. We define the fractional powers of A-A on $\dot{H}$ by
@@
To study the invariant measure, we follow [15] and introduce the spaces corresponding to $\alpha\in\mathbb{R}$:
$\dot{H}^{-1}_{\alpha}:=\left\{v\in\dot{H}^{-1}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}\ \text{and}\ H_{\alpha}:=\left\{v\in\dot{H}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}.$ (2.1)
@@
$e^{-x}\leq Cx^{-\alpha}\quad\text{and}\quad 1-e^{-x}\leq Cx^{\hat{\alpha}},\quad x>0.$ (2.2)
@@
$\left\|(-A)^{\gamma}e^{-A^{2}t}\hat{u}\right\|_{L^{2}}\leq C\left(t^{-\frac{\gamma}{2}}e^{-\frac{\lambda_{1}^{2}}{2}t}\left\|\hat{u}\right\|_{L^{2}}+g(\gamma)\left|\left\langle\hat{u},\phi_{0}\right\rangle_{L^{2}}\right|\right),\quad t>0,\$(2.3)
$\left\|(-A)^{\gamma_{1}}e^{-A^{2}t}\left(I-e^{-A^{2}s}\right)\hat{u}\right\|_{L^{2}}\leq C\left(t^{-\frac{\gamma_{1}-\gamma_{2}+2\gamma_{3}}{2}}s^{\gamma_{3}}e^{-\frac{\lambda_{1}^{2}}{2}t}\left\|(-A)^{\gamma_{2}}\hat{u}\right\|_{L^{2}}+g(\gamma_{1})\left|\left\langle\hat{u},\phi_{0}\right\rangle_{L^{2}}\right|\right),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad s,t>0,$ (2.4)
@@
We next state the assumptions on the nonlinear term $f$, the initial value $u_{0}$, and the noise $W(t,x)$.
Assumption 2.2.
Suppose that $f$ satisfies the following two conditions:
@@
(ii) $L_{f}<\lambda_{1}$, where $L_{f}=-\sup_{x\in\mathcal{O}}f^{\prime}(x)$.
The above assumption ensures that $f$ satisfies
$(y-x)\bigl(f(x)-f(y)\bigr)$ $\leq L_{f}\,|x-y|^{2},\qquad\ \qquad\quad\ x,y\in\mathbb{R},$ (2.5)
$|f(x)-f(y)|$ $\leq C_{f}\bigl(1+x^{2}+y^{2}\bigr)|x-y|,\quad x,y\in\mathbb{R},$ (2.6)
$(1-\epsilon_{0})\left\langle A\hat{u},\hat{u}\right\rangle_{L^{2}}+\left\langle F(\hat{u}-\hat{v})-F(\hat{v}),\hat{u}\right\rangle_{L^{2}}$ $\leq-\epsilon_{0}\|\hat{u}\|_{L^{2}},\qquad\qquad\ \qquad\hat{u},\hat{v}\in L^{6}(\mathcal{O}),$ (2.7)
@@
Assumption 2.3.
Suppose that the initial condition $u_{0}\in H^{1}.$
Assumption 2.4.
Suppose that the noise $W(t,\cdot)$ is a cylindrical Wiener process on $\dot{H}$, which can be represented by the formal series
@@
The concepts of the invariant measure and the Feller property are introduced below, following [24] for details. For a given Banach space $\mathcal{X}$, we denote by $B_{b}(\mathcal{X})$ (resp. $C_{b}(\mathcal{X})$) the Banach space of all bounded, Borel-measurable functions (resp. bounded continuous functions) on $\mathcal{X}$. Furthermore, let $\mathcal{L}(\mathcal{X})$ denote the space of all bounded linear operators on $\mathcal{X}$. We define the transition semigroup $P_{t}:[0,\infty)\rightarrow\mathcal{L}\left(B_{b}(\mathcal{X})\right)$ by
@@
where $u(u_{0};t,\cdot)$ is the solution of the SCHE (1.1) with initial condition $u(0,\cdot)=u_{0}(\cdot)$. It can be verified that $\left\{P_{t}\right\}_{t\geq 0}$ is a Markov semigroup.
@@
In Definition 2.4, we define the invariant measure for the transition semigroup $P_{t}$, which is uniquely determined by the SCHE (1.1). Therefore, in the following, the invariant measure for the semigroup, the SCHE, or the solution will refer to the same object and be used interchangeably.
3.  Invariant measure for the stochastic Cahn–Hilliard equation
This section is devoted to the investigation of the invariant measure for the SCHE (1.1). In [15], the authors established the existence and uniqueness of the invariant measure for (1.1) on the space $\dot{H}^{-1}_{\alpha}$. In this work, we extend these results to the space $H_{\alpha}$. Employing similar arguments to those in [15], we directly conclude that the SCHE (1.1) admits a unique mild solution $u$ given by
$u(t,x)=$ $\int_{\mathcal{O}}G_{t}(x,y)u_{0}\left(y\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta G_{t-s}(x,y)f\left(u\left(s,y\right)\right)\mathrm{d}y\mathrm{d}s+\sigma\sum_{j=1}^{\infty}\int_{0}^{t}\int_{\mathcal{O}}G_{t-s}(x,y)\phi_{j}(y)\mathrm{d}y\mathrm{d}\beta_{j}(s),$ (3.1)
@@ 3.  Invariant measure for the stochastic Cahn–Hilliard equation
$G_{t}(x,y):=\sum_{j=0}^{\infty}e^{-\lambda_{j}^{2}t}\phi_{j}(x)\phi_{j}(y),\quad t\in[0,\infty),\ x,\ y\in\mathcal{O}.$ (3.2)
@@
$\int_{\mathcal{O}}|G_{t}(x,y)-G_{t}(z,y)|^{2}\mathrm{d}y\leq e^{-\lambda_{1}t}\sum_{j=0}^{\infty}\lambda_{j}e^{-\lambda_{j}^{2}t}\left|x-z\right|^{2},\quad t\geq 0,\ x,z\in\mathcal{O},$ (3.3)
$\int_{\mathcal{O}}|\Delta G_{t}(x,y)-\Delta G_{t}(z,y)|\mathrm{d}y\leq Ce^{-\frac{\lambda_{1}}{2}t}\left|x-z\right|^{\alpha}\sqrt{\sum_{j=0}^{\infty}\lambda_{j}^{2+\alpha}e^{-\lambda_{j}^{2}t}},\quad t\geq 0,\ x,z\in\mathcal{O}.$ (3.4)
For the subsequent analysis, we rewrite (1.1) as
$\left\{\begin{aligned} &\mathrm{d}U_{t}+A^{2}U_{t}\mathrm{d}t=AF\left(U_{t}\right)\mathrm{d}t+\sigma\mathrm{d}W_{t},\quad t\in(0,\infty),\\ &U_{0}=u_{0}\in H.\end{aligned}\right$. (3.5)
Eq. (3.5) possesses a unique mild solution given by
@@
3.Key estimates for two auxiliary equations
@@
$\left\{\begin{aligned} &\mathrm{d}V_{t}+A^{2}V_{t}\mathrm{d}t=AF\left(V_{t}+\mathscr{O}_{t}\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &V_{0}=u_{0}.\end{aligned}\right$. (3.6)
@@
$\left\|v\right\|_{L^{6}}\leq C\left(\left\|v\right\|_{L^{2}}^{2}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}^{2}+\left\|Av\right\|_{L^{2}}^{2}\right)^{\frac{1}{12}}\left\|v\right\|_{L^{2}}^{\frac{5}{6}}\leq C\left(\left\|Av\right\|_{L^{2}}^{\frac{1}{6}}\left\|v\right\|_{L^{2}}^{\frac{5}{6}}+\left\|v\right\|_{L^{2}}\right),\quad v\in H^{2},$ (3.7)
@@
$\left\|v\right\|_{L^{\infty}}\leq C\left(\left\|v\right\|_{L^{2}}^{2}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}^{2}\right)^{\frac{1}{4}}\left\|v\right\|_{L^{2}}^{\frac{1}{2}}\leq C\left(\left\|v\right\|_{L^{2}}+\big\|(-A)^{\frac{1}{2}}v\big\|_{L^{2}}\right),\quad v\in H^{1}.$ (3.8)
In the next proposition, we present a general result. In particular, when $Z_{t}=\mathscr{O}_{t}$ and $y_{0}=u_{0}$ (3.9), we obtain the estimate for (3.6).
@@
$\left\{\begin{aligned} &\mathrm{d}Y_{t}+A^{2}Y_{t}\mathrm{d}t=AF\left(Y_{t}+Z_{t}\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &Y_{0}=y_{0},\end{aligned}\right$. (3.9)
@@
$\|Y_{t}\|^{2}_{L^{2}}+\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\|AY_{s}\|^{2}_{L^{2}}\mathrm{d}s\leq C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{2}+K_{1}(t)\right).$ (3.11)
Taking the inner product of (3.9) with $Y_{t}$ yields
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|Y_{t}\right\|^{2}_{L^{2}}\right)=$ $-\left\|AY_{t}\right\|^{2}_{L^{2}}+\left\langle F(Y_{t}+Z_{t}),AY_{t}\right\rangle_{L^{2}}.$ (3.12)
@@
$\left\langle F(Y_{t}+Z_{t}),AY_{t}\right\rangle_{L^{2}}=$ $\int_{\mathcal{O}}\left(a_{0}\left(y(t,x)+z(t,x)\right)^{3}+a_{1}\left(y(t,x)+z(t,x)\right)^{2}+a_{2}\left(y(t,x)+z(t,x)\right)+a_{3}\right)$
$\qquad\qquad\qquad\times\partial_{xx}y(t,x)\mathrm{d}x$